# Synthetic Augmentation in Imbalanced Learning: When It Helps, When It Hurts, and How Much to Add

### v1

@@ 3 Excess Risk Analysis for Imbalanced Classification
In this section, we establish a fundamental limitation on the balanced excess risk when training on imbalanced data augmented with synthetic minority samples. The key obstruction is a distribution bias induced by (i) the effective class weighting in the augmented training population, which typically differs from the target balanced design $(\pi_{0}=1/2)$, and (ii) mismatch between the synthetic distribution $\mathcal{P}_{\mathrm{syn}}$ and the true minority distribution $\mathcal{P}_{1}$. We begin by making these two effects explicit through a decomposition of the synthetic population risk, and then use it to characterize $\mathcal{R}(\hat{\bm{\theta}})-\mathcal{R}(\bm{\theta}^{*})$.
@@
$\tilde{\mathcal{R}}(\bm{\theta})=$ $\,\mathcal{R}(\bm{\theta})+\Big(\pi_{0}-\tfrac{1}{2}\Big)\Big[\mathbb{E}_{\mathcal{P}_{0}}\ell(\bm{\theta};\bm{x},0)-\mathbb{E}_{\mathcal{P}_{1}}\ell(\bm{\theta};\bm{x},1)\Big]+\tilde{\pi}\Big[\mathbb{E}_{\mathcal{P}_{\text{syn}}}\ell(\bm{\theta};\bm{x},1)-\mathbb{E}_{\mathcal{P}_{1}}\ell(\bm{\theta};\bm{x},1)\Big].$ (2)
@@
$\phi(\bm{\theta})$ $:=\mathbb{E}_{\mathcal{P}_{0}}\ell(\bm{\theta};\bm{x},0)-\mathbb{E}_{\mathcal{P}_{1}}\ell(\bm{\theta};\bm{x},1),$ (3)
$\psi(\bm{\theta})$ $:=\mathbb{E}_{\mathcal{P}_{\text{syn}}}\ell(\bm{\theta};\bm{x},1)-\mathbb{E}_{\mathcal{P}_{1}}\ell(\bm{\theta};\bm{x},1),$ (4)
@@
Here, $\big(\pi_{0}-\tfrac{1}{2}\big)\phi(\bm{\theta})+\tilde{\pi}\psi(\bm{\theta})$ is the risk-level distribution bias induced by class-proportion mismatch and generator mismatch. Since our excess-risk analysis is driven by first-order optimality conditions, we will work with the gradient of this bias term. Accordingly, we define the corresponding first-order bias vector
$\bm{b}(\bm{\theta}):=\left(\pi_{0}-\tfrac{1}{2}\right)\nabla\phi(\bm{\theta})+\tilde{\pi}\nabla\psi(\bm{\theta}),$ (6)
which is exactly the $\bm{\theta}$-gradient of the bias component in (5).
@@
Assume the parameter space is bounded, $\bm{\theta}\in\{\bm{\theta}:\|\bm{\theta}\|\leq B\}$. Suppose the covariates $\bm{x}$ have bounded support and the loss $\ell(\bm{\theta};\bm{x},y)$ is continuous in $\bm{\theta}$ and bounded on bounded $(\bm{\theta},\bm{x})$. Assume $\mathcal{R}(\bm{\theta})$ is continuous and $\tilde{\mathcal{R}}(\bm{\theta})$ admits a unique minimizer. Further suppose that $\nabla\tilde{\mathcal{R}}(\bm{\theta})$ is Lipschitz at its minimizer $\tilde{\bm{\theta}}$, i.e.,
$\|\nabla\tilde{\mathcal{R}}(\bm{\theta})-\nabla\tilde{\mathcal{R}}(\tilde{\bm{\theta}})\|\leq M\|\bm{\theta}-\tilde{\bm{\theta}}\|$,
and that $\mathcal{R}(\bm{\theta})$ is (locally) strongly convex at its minimizer $\bm{\theta}^{*}$:
$\mathcal{R}(\bm{\theta})\geq\mathcal{R}(\bm{\theta}^{*})+\frac{\mu}{2}\|\bm{\theta}-\bm{\theta}^{*}\|^{2}.$
Then the balanced excess risk satisfies
@@
Theorem 1 shows that the leading lower bound is governed by the squared norm of the first-order bias vector
$\bm{b}(\bm{\theta}^{*}):=\left(\pi_{0}-\tfrac{1}{2}\right)\nabla\phi(\bm{\theta}^{*})+\tilde{\pi}\nabla\psi(\bm{\theta}^{*}).$
The term proportional to $\pi_{0}-\tfrac{1}{2}$ reflects the class-proportion mismatch in the augmented population, whereas the term proportional to $\tilde{\pi}$ captures generator mismatch through $\mathcal{P}_{\mathrm{syn}}-\mathcal{P}_{1}$. In particular, a necessary condition for balanced-risk consistency, i.e., $\mathcal{R}(\hat{\bm{\theta}})-\mathcal{R}(\bm{\theta}^{*})\to 0$, is $\|\bm{b}(\bm{\theta}^{*})\|\to 0$.
@@
    The balanced risk $\mathcal{R}$ is differentiable and satisfies at its minimizer $\bm{\theta}^{*}$ for a constant $\mu>0$ that
    $\mathcal{R}(\bm{\theta}^{*})\geq\mathcal{R}(\bm{\theta})+\nabla\mathcal{R}(\bm{\theta})^{\top}(\bm{\theta}^{*}-\bm{\theta})+\frac{\mu}{2}\|\bm{\theta}^{*}-\bm{\theta}\|^{2};$
  2. (A2)
    The population risk $\mathcal{R}$ satisfies $\mathcal{R}(\bm{\theta})-\mathcal{R}(\bm{\theta}^{*})\geq L\|\bm{\theta}-\bm{\theta}^{*}\|^{2}$ at the minimizer $\bm{\theta}^{*}$ for some constant $L>0$;
@@
    A finite gradient of total bias:
    $\left\|\left(\pi_{0}-1/2\right)\nabla^{2}\phi(\bm{\theta})+\tilde{\pi}\nabla^{2}\psi(\bm{\theta})\right\|\leq C<\infty.$
@@
Our presentation focuses on the balanced objective $(\rho=\tfrac{1}{2})$ for clarity, but all results extend directly to user-specified class weights. Fix $\rho\in(0,1)$ as the weight on the majority class and define
$\mathcal{R}^{\rho}(\bm{\theta})=\rho\,\mathbb{E}_{\mathcal{P}_{0}}\ell(\bm{\theta};\bm{x},0)+(1-\rho)\,\mathbb{E}_{\mathcal{P}_{1}}\ell(\bm{\theta};\bm{x},1).$
Then the synthetic population risk admits the decomposition
$\tilde{\mathcal{R}}(\bm{\theta})=\mathcal{R}^{\rho}(\bm{\theta})+(\pi_{0}-\rho)\,\phi(\bm{\theta})+\tilde{\pi}\,\psi(\bm{\theta}),$
so the corresponding first-order bias vector becomes
$\bm{b}_{\rho}(\bm{\theta}):=(\pi_{0}-\rho)\,\nabla\phi(\bm{\theta})+\tilde{\pi}\,\nabla\psi(\bm{\theta}).$
Consequently, Theorems 1 and 2 hold verbatim after replacing $\bm{b}(\bm{\theta})$ by $\bm{b}_{\rho}(\bm{\theta})$ (and replacing $\tfrac{1}{2}$ by $\rho$ wherever it appears). Moreover, consistency under the $\rho$-weighted objective requires matching the target proportion $\pi_{0}\to\rho$, i.e.,
$\tilde{n}=\tfrac{1-\rho}{\rho}n_{0}-n_{1}+o(n_{0}),$
and, as in the $\rho=\tfrac{1}{2}$ case, lower-order deviations of $\tilde{n}$ around this target can affect rates.
@@ 4.1 With Ideal Synthetic Sample Generators
Assume the conditions of Theorem 2. Suppose moreover that the curvature contribution from the synthetic mismatch is controlled in the sense that
$\|\nabla^{2}\psi(\bm{\theta}^{*})\|\leq\lambda_{\min}\left(\nabla^{2}\mathcal{R}(\bm{\theta}^{*})\right).$
If the generator is ideal, satisfying $\|\nabla\psi(\bm{\theta}^{*})\|\ll n_{0}^{-1/2}$, then for any synthetic size
$\tilde{n}\in\mathscr{C}:=\left\{\tilde{n}:\ \tilde{n}=n_{0}-n_{1}+o(\sqrt{n_{0}})\right\},$
we have,
$\mathcal{R}(\hat{\bm{\theta}})-\mathcal{R}(\bm{\theta}^{*})=\Theta_{P}\left(n_{0}^{-1}\right).$
@@ 4.2 With Realistic Synthetic Sample Generators
We next consider a more realistic regime in which the synthetic generator is statistically consistent, but its residual mismatch decays more slowly than the parametric rate:
$n_{0}^{-1/2}\ll\|\nabla\psi(\bm{\theta}^{*})\|\ll 1.$
Our goal is to understand how the interaction between the intrinsic majority–minority imbalance and the generator-induced bias determines the appropriate order of the synthetic sample size $\tilde{n}$. To build intuition, we consider the following toy example.
@@
Assume the generator is consistent with $\|\nabla\psi(\bm{\theta}^{*})\|\asymp(\log n_{1})^{-1/2}$, so $\|\nabla\psi(\bm{\theta}^{*})\|\gg n_{0}^{-1/2}$. Write the leading terms in Theorem 2 as
$\mathcal{R}(\hat{\bm{\theta}})-\mathcal{R}(\bm{\theta}^{*})=T_{1}+T_{2}+T_{3}+R,$
where $T_{1}\asymp\|\bm{b}(\bm{\theta}^{*})\|^{2}$, $T_{2}=\Theta_{P}(n^{-1/2}\|\bm{b}(\bm{\theta}^{*})\|)$, and $T_{3}=\Theta_{P}(n^{-1})$, with $n=n_{0}+n_{1}+\tilde{n}$.
Since consistency requires $\tilde{n}\approx n_{0}-n_{1}$, let $\tilde{n}=(n_{0}-n_{1})(1+\delta_{n})$ with $\delta_{n}\to 0$. If $\nabla\psi(\bm{\theta}^{*})$ is collinear with $\nabla\phi(\bm{\theta}^{*})$, i.e., $\nabla\psi(\bm{\theta}^{*})=s(\log n_{1})^{-1/2}\nabla\phi(\bm{\theta}^{*})$ for $s\in\{\pm 1\}$, then one can choose $\delta_{n}=\Theta((\log n_{1})^{-1/2})$ so that $\bm{b}(\bm{\theta}^{*})=0$. Consequently, $T_{1}=T_{2}=0$ and $\mathcal{R}(\hat{\bm{\theta}})-\mathcal{R}(\bm{\theta}^{*})=\Theta_{P}(n_{0}^{-1})$.
In contrast, the naive balancing choice $\tilde{n}=n_{0}-n_{1}$ leaves $\|\bm{b}(\bm{\theta}^{*})\|=\Theta((\log n_{1})^{-1/2})$, yielding a slower rate $\mathcal{R}(\hat{\bm{\theta}})-\mathcal{R}(\bm{\theta}^{*})=\Theta_{P}((\log n_{1})^{-1})$.
@@
Suppose the assumptions of Theorem 2 hold and $\|\nabla^{2}\psi(\bm{\theta}^{*})\|<\infty$. Assume the generator is realistic in the sense that $\|\nabla\psi(\bm{\theta}^{*})\|\gg n_{0}^{-1/2}$, and that the alignment condition
$\sin\angle\left(\nabla\phi(\bm{\theta}^{*}),\nabla\psi(\bm{\theta}^{*})\right)\lesssim\frac{1}{\|\nabla\psi(\bm{\theta}^{*})\|\sqrt{n_{0}}}$
holds. Then:
  • If we choose
    $\tilde{n}=\frac{n_{0}-n_{1}+O_{P}(\sqrt{n_{0}})}{1-2\|\nabla\psi(\bm{\theta}^{*})\|/\|\nabla\phi(\bm{\theta}^{*})\|+O_{P}(1/\sqrt{n_{0}})},$
    then $\mathcal{R}(\hat{\bm{\theta}})-\mathcal{R}(\bm{\theta}^{*})=\Theta_{P}(n_{0}^{-1})$.
  • If we choose the naive balancing rule $\tilde{n}=n_{0}-n_{1}$, then $\mathcal{R}(\hat{\bm{\theta}})-\mathcal{R}(\bm{\theta}^{*})=\Theta_{P}(\|\nabla\psi(\bm{\theta}^{*})\|^{2})\gg\Theta_{P}(n_{0}^{-1})$.
Theorem 4 formalizes a bias-cancellation phenomenon. When the generator mismatch $\nabla\psi(\bm{\theta}^{*})$ is sufficiently aligned with the intrinsic shift $\nabla\phi(\bm{\theta}^{*})$, a small refinement of $\tilde{n}$ around $n_{0}-n_{1}$ cancels the leading distribution bias and restores the parametric rate $\Theta_{P}(n_{0}^{-1})$. In contrast, the naive balancing rule $\tilde{n}=n_{0}-n_{1}$ leaves a residual bias of order $\|\nabla\psi(\bm{\theta}^{*})\|$, resulting in a slower excess-risk rate.

### v2

@@ 3 Excess Risk Analysis for Imbalanced Classification
In this section, we establish a fundamental limitation on the balanced excess risk when training on imbalanced data augmented with synthetic minority samples. The key obstruction is a distribution bias induced by (i) the effective class weighting in the augmented training population and (ii) the mismatch between the synthetic distribution $\mathcal{P}_{\mathrm{syn}}$ and the true minority distribution $\mathcal{P}_{1}$. We begin by making these two effects explicit through a decomposition of the synthetic population risk, and then use it to characterize $\mathcal{R}(\hat{\bm{\theta}})-\mathcal{R}(\bm{\theta}^{*})$.
@@
$\tilde{\mathcal{R}}(\bm{\theta})$ $=\,\mathcal{R}(\bm{\theta})$
$+\Big(\pi_{0}-\tfrac{1}{2}\Big)\Big[\mathbb{E}_{\mathcal{P}_{0}}\ell(\bm{\theta};\bm{x},0)-\mathbb{E}_{\mathcal{P}_{1}}\ell(\bm{\theta};\bm{x},1)\Big]+\tilde{\pi}\Big[\mathbb{E}_{\mathcal{P}_{\text{syn}}}\ell(\bm{\theta};\bm{x},1)-\mathbb{E}_{\mathcal{P}_{1}}\ell(\bm{\theta};\bm{x},1)\Big].$ (2)
@@
$\phi(\bm{\theta}):=\mathbb{E}_{\mathcal{P}_{0}}\ell(\bm{\theta};\bm{x},0)-\mathbb{E}_{\mathcal{P}_{1}}\ell(\bm{\theta};\bm{x},1),\qquad\psi(\bm{\theta}):=\mathbb{E}_{\mathcal{P}_{\text{syn}}}\ell(\bm{\theta};\bm{x},1)-\mathbb{E}_{\mathcal{P}_{1}}\ell(\bm{\theta};\bm{x},1),$ (3)
@@
Here, $\big(\pi_{0}-\tfrac{1}{2}\big)\phi(\bm{\theta})+\tilde{\pi}\psi(\bm{\theta})$ is the risk-level distribution bias induced by class-proportion mismatch and generator mismatch. Since our excess-risk analysis is driven by first-order optimality conditions, we will work with the gradient of this bias term. Accordingly, we define the corresponding first-order bias vector $\bm{b}(\bm{\theta}):=\left(\pi_{0}-\tfrac{1}{2}\right)\nabla\phi(\bm{\theta})+\tilde{\pi}\nabla\psi(\bm{\theta}),$ which is exactly the $\bm{\theta}$-gradient of the bias component in (4).
@@
Assume the parameter space is bounded, $\bm{\theta}\in\{\bm{\theta}:\|\bm{\theta}\|<B\}$. Suppose the covariates $\bm{x}$ have bounded support and the loss $\ell(\bm{\theta};\bm{x},y)$ is continuous and differentiable in $\bm{\theta}$, and bounded on bounded $(\bm{\theta},\bm{x})$. Assume $\mathcal{R}(\bm{\theta})$ is continuous and $\tilde{\mathcal{R}}(\bm{\theta})$ admits a unique minimizer. Further suppose that $\nabla\tilde{\mathcal{R}}(\bm{\theta})$ is Lipschitz at its minimizer $\tilde{\bm{\theta}}$, i.e., $\|\nabla\tilde{\mathcal{R}}(\bm{\theta})-\nabla\tilde{\mathcal{R}}(\tilde{\bm{\theta}})\|\leq M\|\bm{\theta}-\tilde{\bm{\theta}}\|$, and $\mathcal{R}(\bm{\theta})$ satisfies at its minimizer $\bm{\theta}^{*}$ that for any $\theta$: $\mathcal{R}(\bm{\theta})\geq\mathcal{R}(\bm{\theta}^{*})+\frac{\mu}{2}\|\bm{\theta}-\bm{\theta}^{*}\|^{2}.$ Then the balanced excess risk satisfies
@@
Theorem 1 shows that the leading lower bound is governed by the squared norm of the first-order bias vector $\bm{b}(\bm{\theta}^{*}):=\left(\pi_{0}-\tfrac{1}{2}\right)\nabla\phi(\bm{\theta}^{*})+\tilde{\pi}\nabla\psi(\bm{\theta}^{*}).$ The term proportional to $\pi_{0}-\tfrac{1}{2}$ reflects the class-proportion mismatch in the augmented population, whereas the term proportional to $\tilde{\pi}$ captures generator mismatch via the discrepancy between $\mathcal{P}_{\mathrm{syn}}$ and $\mathcal{P}_{1}$. In particular, a necessary condition for balanced-risk consistency, i.e., $\mathcal{R}(\hat{\bm{\theta}})-\mathcal{R}(\bm{\theta}^{*})\to 0$, is $\|\bm{b}(\bm{\theta}^{*})\|\to 0$.
@@
    The balanced risk $\mathcal{R}$ is differentiable and satisfies at its minimizer $\bm{\theta}^{*}$ for a constant $\mu>0$ that $\mathcal{R}(\bm{\theta}^{*})\geq\mathcal{R}(\bm{\theta})+\nabla\mathcal{R}(\bm{\theta})^{\top}(\bm{\theta}^{*}-\bm{\theta})+\frac{\mu}{2}\|\bm{\theta}^{*}-\bm{\theta}\|^{2}$. Also, it satisfies $\mathcal{R}(\bm{\theta})-\mathcal{R}(\bm{\theta}^{*})\geq L\|\bm{\theta}-\bm{\theta}^{*}\|^{2}$ at the minimizer $\bm{\theta}^{*}$ for some constant $L>0$;
@@
    A finite gradient of total bias: $\left\|\left(\pi_{0}-1/2\right)\nabla^{2}\phi(\bm{\theta})+\tilde{\pi}\nabla^{2}\psi(\bm{\theta})\right\|\leq C<\infty.$
@@
Our presentation focuses on the balanced objective $(\rho=\tfrac{1}{2})$ for clarity, but all results extend directly to user-specified class weights. Fix $\rho\in(0,1)$ as the weight on the majority class and define $\mathcal{R}^{\rho}(\bm{\theta})=\rho\,\mathbb{E}_{\mathcal{P}_{0}}\ell(\bm{\theta};\bm{x},0)+(1-\rho)\,\mathbb{E}_{\mathcal{P}_{1}}\ell(\bm{\theta};\bm{x},1).$ Then the synthetic population risk admits the decomposition $\tilde{\mathcal{R}}(\bm{\theta})=\mathcal{R}^{\rho}(\bm{\theta})+(\pi_{0}-\rho)\,\phi(\bm{\theta})+\tilde{\pi}\,\psi(\bm{\theta}),$ so the corresponding first-order bias vector becomes $\bm{b}_{\rho}(\bm{\theta}):=(\pi_{0}-\rho)\,\nabla\phi(\bm{\theta})+\tilde{\pi}\,\nabla\psi(\bm{\theta}).$ Consequently, Theorems 1 and 2 hold after replacing $\bm{b}(\bm{\theta})$ by $\bm{b}_{\rho}(\bm{\theta})$ (and replacing $\tfrac{1}{2}$ by $\rho$ wherever it appears). Moreover, consistency under the $\rho$-weighted objective requires matching the target proportion $\pi_{0}\to\rho$, i.e., $\tilde{n}=\tfrac{1-\rho}{\rho}n_{0}-n_{1}+o(n_{0}),$ and, as in the $\rho=\tfrac{1}{2}$ case, lower-order deviations of $\tilde{n}$ around this target can affect rates.
@@ 4.1 With Ideal Synthetic Sample Generators
Assume the conditions of Theorem 2. Suppose that the curvature contribution from the synthetic mismatch is controlled in the sense that $\|\nabla^{2}\psi(\bm{\theta}^{*})\|\leq\lambda_{\min}\left(\nabla^{2}\mathcal{R}(\bm{\theta}^{*})\right).$ If the generator is ideal, satisfying $\|\nabla\psi(\bm{\theta}^{*})\|\ll n_{0}^{-1/2}$, then for any synthetic size $\tilde{n}\in\mathscr{C}:=\left\{\tilde{n}:\ \tilde{n}=n_{0}-n_{1}+o(\sqrt{n_{0}})\right\},$ we have $\mathcal{R}(\hat{\bm{\theta}})-\mathcal{R}(\bm{\theta}^{*})=\Theta_{P}\left(n_{0}^{-1}\right).$
@@ 4.2 With Realistic Synthetic Sample Generators
We next consider a more realistic regime in which the synthetic generator is statistically consistent, but its residual mismatch decays more slowly than the parametric rate: $n_{0}^{-1/2}\ll\|\nabla\psi(\bm{\theta}^{*})\|\ll 1.$ Our goal is to understand how the interaction between the intrinsic majority–minority imbalance and the generator-induced bias determines the appropriate order of the synthetic sample size $\tilde{n}$. To build intuition, we consider the following toy example.
@@
Assume the generator is consistent with $\|\nabla\psi(\bm{\theta}^{*})\|\asymp(\log n_{1})^{-1/2}$, so $\|\nabla\psi(\bm{\theta}^{*})\|\gg n_{0}^{-1/2}$. Write the leading terms in Theorem 2 as $\mathcal{R}(\hat{\bm{\theta}})-\mathcal{R}(\bm{\theta}^{*})=T_{1}+T_{2}+T_{3}+R,$ where $T_{1}\asymp\|\bm{b}(\bm{\theta}^{*})\|^{2}$, $T_{2}=\Theta_{P}(n^{-1/2}\|\bm{b}(\bm{\theta}^{*})\|)$, and $T_{3}=\Theta_{P}(n^{-1})$, with $n=n_{0}+n_{1}+\tilde{n}$. Since consistency requires $\tilde{n}\approx n_{0}-n_{1}$, let $\tilde{n}=(n_{0}-n_{1})(1+\delta_{n})$ with $\delta_{n}\to 0$. If $\nabla\psi(\bm{\theta}^{*})$ is collinear with $\nabla\phi(\bm{\theta}^{*})$, i.e., $\nabla\psi(\bm{\theta}^{*})=s(\log n_{1})^{-1/2}\nabla\phi(\bm{\theta}^{*})$ for $s\in\{\pm 1\}$, then one can choose $\delta_{n}=\Theta((\log n_{1})^{-1/2})$ so that $\bm{b}(\bm{\theta}^{*})=0$. Consequently, $T_{1}=T_{2}=0$ and $\mathcal{R}(\hat{\bm{\theta}})-\mathcal{R}(\bm{\theta}^{*})=\Theta_{P}(n_{0}^{-1})$. In contrast, the naive balancing choice $\tilde{n}=n_{0}-n_{1}$ leaves $\|\bm{b}(\bm{\theta}^{*})\|=\Theta((\log n_{1})^{-1/2})$, yielding a slower rate $\mathcal{R}(\hat{\bm{\theta}})-\mathcal{R}(\bm{\theta}^{*})=\Theta_{P}((\log n_{1})^{-1})$.
@@
Suppose the assumptions of Theorem 2 hold and $\|\nabla^{2}\psi(\bm{\theta}^{*})\|<\infty$. Assume the generator is realistic in the sense that $\|\nabla\psi(\bm{\theta}^{*})\|\gg n_{0}^{-1/2}$, and that the alignment condition $\sin\angle\left(\nabla\phi(\bm{\theta}^{*}),\nabla\psi(\bm{\theta}^{*})\right)\lesssim\frac{1}{\|\nabla\psi(\bm{\theta}^{*})\|\sqrt{n_{0}}}$ holds. Then, if we choose $\tilde{n}=\frac{n_{0}-n_{1}+O_{P}(\sqrt{n_{0}})}{1-2\|\nabla\psi(\bm{\theta}^{*})\|/\|\nabla\phi(\bm{\theta}^{*})\|\cos\angle\left(\nabla\phi(\bm{\theta}^{*}),\nabla\psi(\bm{\theta}^{*})\right)+O_{P}(1/\sqrt{n_{0}})},$ we have $\mathcal{R}(\hat{\bm{\theta}})-\mathcal{R}(\bm{\theta}^{*})=\Theta_{P}(n_{0}^{-1})$. If we choose the naive balancing rule $\tilde{n}=n_{0}-n_{1}$, then $\mathcal{R}(\hat{\bm{\theta}})-\mathcal{R}(\bm{\theta}^{*})=\Theta_{P}(\|\nabla\psi(\bm{\theta}^{*})\|^{2})\gg\Theta_{P}(n_{0}^{-1})$.
Theorem 4 formalizes a “bias-cancellation” phenomenon. When the generator mismatch $\nabla\psi(\bm{\theta}^{*})$ is sufficiently aligned with the intrinsic shift $\nabla\phi(\bm{\theta}^{*})$, a small refinement of $\tilde{n}$ around $n_{0}-n_{1}$ cancels the leading distribution bias and restores the parametric rate $\Theta_{P}(n_{0}^{-1})$. In contrast, the naive balancing rule $\tilde{n}=n_{0}-n_{1}$ leaves a residual bias of order $\|\nabla\psi(\bm{\theta}^{*})\|$, resulting in a slower excess-risk rate.